# Benefits and Pitfalls of Reinforcement Learning for Language Model Planning: A Theoretical Perspective

### v1

@@ Appendix E Appendix for Q-Learning
Fix any triple $(i,j,k)$. Consider training sequences whose first two nodes satisfy $u_{1}\in\mathcal{V}$ and $u_{2}=i$. By the definition of the training process, $\mathbb{P}(u_{1}\in\mathcal{V},u_{2}=i)>0$. Under $\epsilon$-exploration uniform over $\mathcal{V}$,
@@
Condition on the event $\{u_{1}\in\mathcal{V},u_{2}=i\}$. Then in the next two decisions,
$\mathbb{P}(u_{4}=j\mid u_{1}\in\mathcal{V},u_{2}=i)\geq\epsilon/|\mathcal{V}|,\qquad\mathbb{P}(u_{5}=k\mid u_{1}\in\mathcal{V},u_{2}=i,\ u_{4}=j)\geq\epsilon/|\mathcal{V}|.$
@@
$\mathbb{P}(u_{2}=i,u_{4}=j,u_{5}=k)\geq p_{0}(\epsilon/|\mathcal{V}|)^{2}>0.$
Each occurrence of $(u_{2},u_{4},u_{5})=(i,j,k)$ triggers one update of $\mathbf{f}(i,j)[k]$. Since each occurrence yields an update of $\mathbf{f}(i,j)[k]$, we obtain
@@
$\ell=\big(\mathbf{f}(u_{2},u_{m})[u_{m+1}]-\delta_{{\bf u}\in\mathcal{P}}\,\delta_{u_{m+1}=u_{2}}-\{\max_{k^{\prime}}\mathbf{f}(u_{2},u_{m+1})[k^{\prime}]\}\big)^{2}.$
Taking the gradient with respect to $\mathbf{f}(u_{2},u_{m})[u_{m+1}]$ and setting the expectation to zero yields, for every triple $(i,j,k)$,
@@
This expression no longer depends on $k$ or $j$.
@@
$\ell=\big(\mathbf{f}(u_{2},u_{m})[u_{m+1}]-(\delta_{u_{m+1}=u_{2}}-\delta_{(u_{m},u_{m+1})\notin\mathcal{E}})-\{\max_{k^{\prime}}\mathbf{f}(u_{2},u_{m+1})[k^{\prime}]\}\big)^{2}.$
Taking the gradient with respect to the active coordinate $\mathbf{f}(u_{2},u_{m})[u_{m+1}]$ gives
$\frac{\partial\ell}{\partial\mathbf{f}(u_{2},u_{m})[u_{m+1}]}=2\big(\mathbf{f}(u_{2},u_{m})[u_{m+1}]-\delta_{u_{m+1}=u_{2}}+\delta_{(u_{m},u_{m+1})\notin\mathcal{E}}-\max_{k^{\prime}}\mathbf{f}(u_{2},u_{m+1})[k^{\prime}]\big).$
@@
$\mathbf{f}(u_{2},u_{m})[u_{m+1}]\leftarrow(1-2\eta)\,\mathbf{f}(u_{2},u_{m})[u_{m+1}]+2\eta\big(\delta_{u_{m+1}=u_{2}}-\delta_{(u_{m},u_{m+1})\notin\mathcal{E}}+\max_{k^{\prime}}\mathbf{f}(u_{2},u_{m+1})[k^{\prime}]\big).$
Renaming $(i,j,k)=(u_{2},u_{m},u_{m+1})$ and writing $S^{(t)}_{i,k}=\max_{k^{\prime}}\mathbf{f}^{(t)}(i,k)[k^{\prime}]$, the recursion is
@@
Let $N^{\mathrm{prop}}_{i,j,k}$ denote the asymptotic proportion of triples $(u_{2},u_{m},u_{m+1})=(i,j,k)$ occurring in the generated sequences at the stable point, under the given sampling method and the persistent exploration condition. Equivalently, $N^{\mathrm{prop}}_{i,j,k}$ is the limiting frequency with which state $j$ transitions to $k$ with target $i$ in the trajectories sampled by the model. By definition $N^{\mathrm{prop}}_{i,j,k}>0$ for all $i,j,k$.
@@ Appendix G Additional Experimental Results
G.1 Experiments on Erdős-Rényi Graphs
@@
During the RL process, the model generates paths for pairs in $D_{\mathrm{RL\text{-}Train}}$ and receives reward signals. The main difference between this setup and that in Section 2 is that the RL training set now contains new pairs that were unseen during SFT ($D_{\mathrm{Test2Train}}$). Therefore, the initial model is not perfect on these new training pairs. Additionally, some pairs from the SFT training set are not used for RL training ($D_{\mathrm{Train2Test}}$), which allows us to measure the extent of forgetting. We consider the same RL algorithms introduced in Section 2: PG and Q-learning, whose training curves are presented in Figures 5 and 6, respectively. All accuracies are evaluated using greedy decoding.
@@
G.2 Experiments on Graph Represented for Blocksworld

### v2

@@ Appendix E Appendix for Q-Learning
Fix any triple $(i,j,k)$. Consider training sequences whose first two nodes satisfy $u_{\text{source}}\in\mathcal{V}$ and $u_{\text{target}}=i$. By the definition of the training process, $\mathbb{P}(u_{\text{source}}\in\mathcal{V},u_{\text{target}}=i)>0$. Under $\epsilon$-exploration uniform over $\mathcal{V}$,
@@
Condition on the event $\{u_{\text{source}}\in\mathcal{V},u_{\text{target}}=i\}$. Then in the next two decisions,
$\mathbb{P}(u_{2}=j\mid u_{\text{source}}\in\mathcal{V},u_{\text{target}}=i)\geq\epsilon/|\mathcal{V}|,\quad\mathbb{P}(u_{3}=k\mid u_{\text{source}}\in\mathcal{V},u_{\text{target}}=i,\ u_{2}=j)\geq\epsilon/|\mathcal{V}|.$
@@
$\mathbb{P}(u_{\text{target}}=i,u_{2}=j,u_{3}=k)\geq p_{0}(\epsilon/|\mathcal{V}|)^{2}>0.$
Each occurrence of $(u_{\text{target}},u_{2},u_{3})=(i,j,k)$ triggers one update of $\mathbf{f}(i,j)[k]$. Since each occurrence yields an update of $\mathbf{f}(i,j)[k]$, we obtain
@@
$\ell=\big(\mathbf{f}(u_{\text{target}},u_{m})[u_{m+1}]-\delta_{{\bf u}\in\mathcal{P}}\,\delta_{u_{m+1}=u_{\text{target}}}-\{\max_{k^{\prime}}\mathbf{f}(u_{\text{target}},u_{m+1})[k^{\prime}]\}\big)^{2}.$
Taking the gradient with respect to $\mathbf{f}(u_{\text{target}},u_{m})[u_{m+1}]$ and setting the expectation to zero yields, for every triple $(i,j,k)$,
@@
Here the second equality separates the case $k^{\prime}=i$ from $k^{\prime}\neq i$: if $k^{\prime}=i$, the chain terminates immediately, corresponding to the $\cdots\rightarrow k\rightarrow i$ path, so we take the value $\mathbf{f}(i,k)[i]$ directly. If $k^{\prime}\neq i$, we must expand one step further, leading to $\max_{k^{\prime\prime}}\mathbf{f}(i,k^{\prime})[k^{\prime\prime}]$. The last equality then observes that including the $k^{\prime}=i$ term does not change the maximization, since it is already dominated by the expansion over $k^{\prime}\neq i$. Thus the final expression simply enumerates all possible terms with the restriction $k^{\prime}\neq i$. This expression no longer depends on $k$ or $j$.
@@
$\ell=\big(\mathbf{f}(u_{\text{target}},u_{m})[u_{m+1}]-(\delta_{u_{m+1}=u_{\text{target}}}-\delta_{(u_{m},u_{m+1})\notin\mathcal{E}})-\{\max_{k^{\prime}}\mathbf{f}(u_{\text{target}},u_{m+1})[k^{\prime}]\}\big)^{2}.$
Taking the gradient with respect to the active coordinate $\mathbf{f}(u_{\text{target}},u_{m})[u_{m+1}]$ gives
$\frac{\partial\ell}{\partial\mathbf{f}(u_{\text{target}},u_{m})[u_{m+1}]}=2\big(\mathbf{f}(u_{\text{target}},u_{m})[u_{m+1}]-\delta_{u_{m+1}=u_{\text{target}}}+\delta_{(u_{m},u_{m+1})\notin\mathcal{E}}-\max_{k^{\prime}}\mathbf{f}(u_{\text{target}},u_{m+1})[k^{\prime}]\big).$
@@
$\mathbf{f}(u_{\text{target}},u_{m})[u_{m+1}]\leftarrow(1-2\eta)\,\mathbf{f}(u_{\text{target}},u_{m})[u_{m+1}]+2\eta\big(\delta_{u_{m+1}=u_{\text{target}}}-\delta_{(u_{m},u_{m+1})\notin\mathcal{E}}+\max_{k^{\prime}}\mathbf{f}(u_{\text{target}},u_{m+1})[k^{\prime}]\big).$
Renaming $(i,j,k)=(u_{\text{target}},u_{m},u_{m+1})$ and writing $S^{(t)}_{i,k}=\max_{k^{\prime}}\mathbf{f}^{(t)}(i,k)[k^{\prime}]$, the recursion is
@@
Let $N^{\mathrm{prop}}_{i,j,k}$ denote the asymptotic proportion of triples $(u_{\text{target}},u_{m},u_{m+1})=(i,j,k)$ occurring in the generated sequences at the stable point, under the given sampling method and the persistent exploration condition. Equivalently, $N^{\mathrm{prop}}_{i,j,k}$ is the limiting frequency with which state $j$ transitions to $k$ with target $i$ in the trajectories sampled by the model. By definition $N^{\mathrm{prop}}_{i,j,k}>0$ for all $i,j,k$.
@@ Appendix G Additional Experimental Results
(a) Attention weights in SFT after different training iterations
(b) Attention weights in PG ($\lambda=0$) after different training iterations
(c) Attention weights in Q-learning (Process Reward) after different training iterations
Figure 5: Empirical validation that the trained one-layer one-head transformer acts as a function of the target node and the current node. The visualization of attention maps across SFT, PG, and Q-learning training shows a consistent, strong focus on the target node (token position 1).
G.1 Validation of Learned Attention
This subsection visualizes the evolution of attention maps during training for SFT, PG, and Q-learning. Our analysis first focuses on a one-layer, one-head transformer at various training steps (Figure 5). For each model, we compute the average attention weight over the SFT training dataset $\mathcal{D}^{\text{SFT}}$.
Specifically, the weight at position $i$ in row $k$ (where $i\leq k$) represents the average attention the model assigns to the $i$-th token when predicting the $(k+1)$-th token, averaged over all paths in $\mathcal{D}^{\text{SFT}}$ longer than $k+1$. Since the underlying graph has 100 sparse nodes, path lengths in $\mathcal{D}^{\text{SFT}}$ are generally short; consequently, we display only the first 8 rows.
Our visualizations reveal that during SFT, the transformer allocates most attention to the target node. Combined with the residual connections, which allow access to the current node’s information, this suggests that the model learns to predict the next node based primarily on the target and current nodes. This empirical finding aligns with the results of Wang et al. (2024b). Another interesting phenomenon in SFT is that the attention weight on the target node quickly peaks and then gradually decreases, while remaining dominant. We hypothesize that this may be due to overfitting on $\mathcal{D}^{\text{SFT}}$, leading the model to develop auxiliary prediction strategies. This overfitting could also explain the decreasing generalization performance of SFT observed in our experiments.
In contrast, for both PG and Q-learning, the attention on the target node increases throughout training. In Q-learning, the final attention weight on the target node exceeds 95%, making it the closest to the conditions outlined in Assumption 3.1.
To further validate our findings, we extend the analysis to a two-layer, one-head transformer. As shown in Figure 6, which displays the attention map averaged over $\mathcal{D}^{\text{SFT}}$, both layers predominantly attend to the target and current nodes. This pattern strongly supports our assumption that the transformer operates as a function of the target and current nodes, confirming the consistency of this behavior across model architectures.
(a) Attention weights in SFT after different training iterations
(b) Attention weights in PG ($\lambda=0$) after different training iterations
(c) Attention weights in Q-learning (Process Reward) after different training iterations
Figure 6: Empirical validation that the trained two-layer one-head transformer acts as a function of the target and current nodes.
G.2 Experiments on Erdős-Rényi Graphs
@@
During the RL process, the model generates paths for pairs in $D_{\mathrm{RL\text{-}Train}}$ and receives reward signals. The main difference between this setup and that in Section 2 is that the RL training set now contains new pairs that were unseen during SFT ($D_{\mathrm{Test2Train}}$). Therefore, the initial model is not perfect on these new training pairs. Additionally, some pairs from the SFT training set are not used for RL training ($D_{\mathrm{Train2Test}}$), which allows us to measure the extent of forgetting. We consider the same RL algorithms introduced in Section 2: PG and Q-learning, whose training curves are presented in Figures 7 and 8, respectively. All accuracies are evaluated using greedy decoding.
@@
G.3 Experiments on Graph Represented for Blocksworld